OpenStudy (anonymous):

Evaluate: Log_{5} 16 / Log_{5} 4 (no calculator)

OpenStudy (anonymous):

\[2^{-x} = 8\] find exact solution, use common log and two decimal place approximation of each

OpenStudy (radar):

x=(-3)\[2^{-(-3)}=2^{3}=8\]

OpenStudy (radar):

Either I worked it wrong, but that one can be solved by inspection. Now I will think about solving it with logs.

OpenStudy (anonymous):

Yes, please. It is in my precalculs book and they are using logs in this chapter. Thank you.

OpenStudy (radar):

I am very rusty with logs lol. But here is my go at it. -x log2 = log 8 -0.301029x=0.90308998 x=-0.90308998/0.301029=-3.000009923 for 2 decimal place approximation x=-3.00

OpenStudy (radar):

Looks like it still is a -3

OpenStudy (anonymous):

Thank you!